# XOXO: Stealthy Cross-Origin Context Poisoning Attacks against AI Coding Assistants

### v1

@@ 7.1 RQ1: In-Context Code Generation
HumanEval+ MBPP+ Model Attack ASR # Queries ASR # Queries Claude 3.5 Sonnet GCGS 92.00 145 98.42 75 GPT-4o GCGS+P 33.81 378 66.43 161 Artigenz Coder GCGS 88.95 ±3.6 160 ±22 99.25 ±0.68 29 ±3 GCGS+P 92.77 ±1.39 131 ±7 99.88 ±0.28 22 ±3 CodeQwen 1.5 GCGS 87.68 ±0.40 179 ±6 94.84 ±0.93 98 ±5 GCGS+P 91.49 ±2.22 158 ±20 96.09 ±0.57 93 ±7 DeepSeek Coder GCGS 88.36 ±1.75 165 ±10 99.55 ±0.48 25 ±3 GCGS+P 90.73 ±1.63 154 ±9 99.89 ±0.25 20 ±2 Llama 3 GCGS 98.87 ±1.69 54 ±11 96.60 ±0.97 56 ±6 GCGS+P 100.00 ±0.00 39 ±9 97.34 ±1.00 53 ±6
TABLE IV: Naturalness of GCGS and GCGS+P (perplexity-guided) attacks on code generation using HumanEval+ and MBPP+. Results show mean ± std over 5 seeds for open-source models and single runs for closed-source models (limited 5-run analysis in Appendix F).
HumanEval+ MBPP+
Model Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU
Claude 3.5 Sonnet GCGS 1.00 2.86 98.48 1.00 1.82 98.51
GPT-4o GCGS+P 1.91 4.47 97.90 2.68 6.11 96.01
Artigenz Coder GCGS 1.00 ±0.00 1.94 ±0.16 98.70 ±0.07 1.00 ±0.00 1.72 ±0.09 98.33 ±0.09
GCGS+P 2.37 ±0.37 5.67 ±0.91 97.00 ±0.49 1.33 ±0.12 2.35 ±0.22 97.91 ±0.16
CodeQwen 1.5 GCGS 1.00 ±0.00 1.60 ±0.12 97.10 ±0.21 1.00 ±0.00 1.24 ±0.05 99.02 ±0.05
GCGS+P 2.24 ±0.26 5.50 ±0.90 96.25 ±0.48 1.78 ±0.09 3.50 ±0.26 97.61 ±0.18
DeepSeek Coder GCGS 1.00 ±0.00 1.84 ±0.08 98.08 ±0.06 1.00 ±0.00 1.74 ±0.08 98.40 ±0.09
GCGS+P 2.27 ±0.19 5.45 ±0.62 96.98 ±0.45 1.14 ±0.06 2.13 ±0.21 98.16 ±0.11
Llama 3 GCGS 1.00 ±0.00 1.53 ±0.11 98.73 ±0.03 1.00 ±0.00 1.51 ±0.08 98.66 ±0.06
GCGS+P 1.53 ±0.29 3.29 ±0.91 98.12 ±0.33 1.29 ±0.11 2.29 ±0.31 98.18 ±0.19
Attack Performance. As shown in Table III, On all models except GPT-4o, our attacks achieved consistently high success rates: over 87% on HumanEval+ and above 94% on MBPP+. Notably, the perplexity-guided GCGS (GCGS+P) consistently outperforms base GCGS across all open-source models, achieving success-rate improvements of up to 3.82%. The attack is particularly effective against Llama 3, reaching a perfect 100% success rate on HumanEval+ with perplexity-guided GCGS. Note that Claude 3.5 Sonnet demonstrates high vulnerability with a 92% attack success rate on HumanEval+ and 98.42% attack success rate on MBPP+, and only supporting our weaker GCGS variant. The attacks demonstrate practical efficiency, with target models requiring between 20-378 queries per successful attack across both datasets. While GCGS is still successful in attacking GPT-4o, its performance suffers compared to other models. Unfortunately, due to GPT-4o’s closed-source nature, we cannot determine the root cause. One factor is that the prohibitive costs of running GPT-4o forced us to considerably limit the scale of GCGS on GPT-4o compared to other models.
Attack Naturalness. In Table IV, we see the adversarial examples maintain high naturalness across all models, as evidenced by CodeBLEU scores consistently above 96. The base GCGS achieves slightly higher CodeBLEU but at the cost of limited modification scope. In contrast, perplexity-guided GCGS makes more extensive but still natural modifications, affecting more identifiers and positions while maintaining comparable CodeBLEU scores. This suggests that perplexity-guided GCGS finds a better balance between attack effectiveness and naturalness.
{longfbox}
Result 1: GCGS, with or without perplexity guidance, achieves success rates of 92-100% attacking leading frontier models like Claude 3.5 Sonnet and Llama 3.
7.2 RQ2: In-Context Vulnerability Injection
@@
Task Description. To quantify the ability of GCGS+P to inject vulnerabilities while maintaining functional correctness, we evaluate GCGS+P on the CWEval dataset [15], specifically designed to assess both functionality and security of LLM-generated code. Using our identical EvalPlus setup for RQ1 on CWEval’s Python subset, we measured attack success by target LLMs generating code that passes functional tests while failing security tests linked to specific Common Weakness Enumeration (CWE) categories [49]. We additionally evaluate on Codestral 22B v0.1 (outperforms GPT-4o, second after Claude 3.5 Sonnet in Copilot Arena [50]) and Qwen2.5 Coder 32B Instruct (outperforms GPT-4o on EvalPlus). The attack results alongside baseline performance on this dataset are shown in Table V.
TABLE V: pass@1 baseline performance of SoTA LLMs on CWEval, passing both functional & security testcases, alongside performance of GCGS+P (perplexity-guided).
Model pass@1 ASR # Queries # Identifiers # Positions CodeBLEU Artigenz Coder 52.00 58.33 3566 3.43 5.71 97.70 CodeQwen 1.5 32.00 37.50 4358 4.00 8.00 97.57 Codestral 22B v0.1 40.00 66.67 2205 1.00 1.83 99.18 DeepSeek Coder 32.00 75.00 2129 4.50 8.67 97.51 Llama 3 32.00 50.00 5322 1.00 1.75 99.41 Qwen 2.5 Coder 52.00 42.86 4854 4.33 7.67 97.97
We found that GCGS+P successfully triggered 11 unique CWEs across different models, achieving an average attack success rate of 55.06%. The attack required only one identifier change for several models, demonstrating its subtlety. We’ve added three case studies highlighting the most concerning behaviors observed in the tested models in Appendix G.
{longfbox}
Result 2: GCGS can make SoTA LLMs generate vulnerable code with an ASR up to 75% with certain attacks requiring as small as one identifier change.
7.3 RQ3: Code Reasoning
@@
Task Description. To evaluate our ability to attack code reasoning LLMs, we select two security-focused binary classification benchmarks from CodeXGLUE [51]: Defect Detection and Clone Detection, both well-established in the adversarial code transformation literature [28, 29, 52]. The Defect Detection task builds on Devign [53], a dataset of 27,318 real-world C functions extracted from open-source projects and annotated for security vulnerabilities. The Clone Detection task employs BigCloneBench [54, 55], a large-scale dataset with over 1.7 million labeled code pairs spanning from syntactically identical to semantically similar code fragments. We evaluate our attack on three fine-tuned LLMs that achieve SoTA performance on these tasks: CodeBERT [56], GraphCodeBERT [57], and CodeT5+ 110M [58]. While we considered evaluating generative coding models in a zero-shot chat setting, our experiments (shown in the upper part of Table VI) revealed they performed at random chance (50% accuracy) on both binary classification tasks, even with extensive prompt engineering. This poor baseline performance, which aligns with existing findings on LLMs’ limitations in vulnerability detection [59], led us to focus our evaluation on fine-tuned LLMs that show meaningful accuracy on these tasks.
TABLE VI: Performance comparison (accuracy%) of zero-shot generative models against fine-tuned classifier models.
Zero-shot Generation Defect Detection Clone Detection
CodeQwen 1.5 52.33 52.25
DeepSeek Coder 47.36 53.00
Llama 3 44.98 52.25
Fine-tuned Classifiers Defect Detection Clone Detection
CodeBERT 62.03 ±0.88 90.05 ±1.33
GraphCodeBERT 62.95 ±0.62 97.30 ±0.19
CodeT5+ 110M 61.74 ±1.07 84.95 ±2.06
We compare against several leading adversarial attacks that leverage semantics-preserving code transformations: ALERT [28] and MHM [33] (chosen for their prevalence in comparative studies), RNNS [29] (a recent performant approach), and WIR-Random [34] (the most effective non-Java-specific attack from a comprehensive study [35]).
For model training and evaluation, we use different approaches for our two datasets. For Defect Detection, we fine-tune models on the full dataset. For Clone Detection, due to its substantial size, we follow previous literature and use a balanced subset of 90,000 training and 4,000 validation examples to ensure computational feasibility. We sample 400 test examples from Clone Detection to enable multiple evaluations of each attack-model combination. Since all of our baselines also use tree-sitter, we omit test examples where tree-sitter detects parsing errors. For GCGS’s warm-up, we withhold a small subset of the fine-tuning datasets: 1,000 training and 100 validation examples for Defect Detection, and 4,000 training and 200 validation examples for Clone Detection. The one-time computational costs of warm-up in terms of model queries are detailed in Appendix D. To mitigate the effects of randomness during model fine-tuning and attacking, we fine-tune each model five times on five random seeds and run each attack with the same random seed on each fine-tuned model.
Defect Detection. GCGS champions over over existing approaches in both effectiveness and efficiency (see Table VII). GCGS uses up to 42.92% fewer queries than RNNS while delivering consistently higher success rates across all evaluated models. While WIR achieves lower query counts on CodeBERT and GraphCodeBERT, its success rate falls short of GCGS by a considerable margin of up to 28.36%. Warm-up (GCGS+W) further boosts GCGS’s performance. GCGS+W is particularly performant on CodeT5+, where it not only approaches perfect attack success but also dramatically reduces the required queries by 74.01% to just 46 queries on average. Remarkably, GCGS+W achieves this by warming up on just 1,100 examples - a mere 4.47% of the data withheld from fine-tuning.
@@
TABLE VII: Performance of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Defect Detection. Results show mean ± std over 5 seeds. Each best ASR per model is in bold.
@@
CodeBERT GraphCodeBERT CodeT5+ Attack ASR # Queries ASR # Queries ASR # Queries ALERT 62.35 ±5.92 732 ±120 76.87 ±5.00 468 ±106 62.22 ±8.02 784 ±196 MHM 56.48 ±9.14 742 ±111 75.64 ±12.84 479 ±178 82.81 ±1.98 405 ±34 RNNS 73.97 ±6.39 479 ±67 86.51 ±5.11 331 ±61 86.59 ±3.21 355 ±44 WIR 64.82 ±6.65 145 ±11 78.80 ±8.44 125 ±15 74.43 ±2.02 134 ±8 GCGS 93.18 ±5.79 259 ±172 94.11 ±6.13 229 ±155 97.76 ±1.12 177 ±27 GCGS+W 97.17 ±1.90 167 ±45 97.22 ±3.03 147 ±113 99.89 ±0.13 46 ±48
As shown in Table VIII, GCGS outperforms baselines in code naturalness, averaging only 1.84 identifier changes and 9.65 position modifications. Likewise, its CodeBLEU score of 92.94 exceeds WIR’s 86.05. With warm-up, GCGS+W further improves, requiring just 1.05 identifier and 2.61 position changes when attacking CodeT5+.
TABLE VIII: Naturalness of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Defect Detection. Results show mean ± std over 5 seeds. Each best score per model is bold.
CodeBERT GraphCodeBERT CodeT5+ Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU ALERT 3.01 ±0.23 25.42 ±1.85 81.59 ±1.72 2.62 ±0.22 20.04 ±2.34 84.35 ±0.87 2.95 ±0.25 23.68 ±1.22 82.91 ±0.56 MHM 2.74 ±0.30 20.54 ±2.79 84.67 ±1.19 2.59 ±0.21 17.88 ±2.08 86.05 ±0.78 2.75 ±0.16 19.45 ±1.58 84.18 ±0.75 RNNS 3.92 ±0.59 32.45 ±6.23 86.85 ±1.10 2.60 ±0.49 22.43 ±4.66 88.01 ±0.88 2.76 ±0.29 23.53 ±3.59 87.74 ±0.94 WIR 2.64 ±0.17 21.99 ±2.03 85.05 ±0.99 2.22 ±0.26 17.22 ±2.55 86.91 ±0.72 2.40 ±0.20 18.39 ±1.28 86.18 ±0.73 GCGS 2.00 ±0.26 9.96 ±2.33 92.83 ±1.28 1.94 ±0.24 11.51 ±2.55 91.61 ±1.29 1.57 ±0.06 7.48 ±0.53 94.37 ±0.38 GCGS+W 1.49 ±0.18 5.94 ±1.38 95.62 ±0.94 1.45 ±0.26 7.83 ±2.85 94.13 ±2.00 1.05 ±0.03 2.61 ±0.90 97.93 ±0.77

### v2

@@ 6.1 In-Context Code Generation
HumanEval+ MBPP+ CWEval/Python Model Attack ASR # Queries ASR # Queries ASR # Queries Claude 3.5 Sonnet v2 GCGS 92.00 145 98.42 75 40.00 4690 GPT 4.1 GCGS+P 81.82 150 40.69 233 50.00 4144 Codestral 22B GCGS 74.15 ±0.89 273 ±7 98.99 ±0.60 43 ±3 60.30 ± 4.81 3077 ±234 GCGS+P 78.70 ±1.85 263 ±13 99.36 ±0.25 37 ±1 62.58 ± 5.76 2927 ±221 DeepSeek Coder 6.7B GCGS 88.36 ±1.75 165 ±10 99.55 ±0.48 25 ±3 64.44 ± 9.30 3128 ±218 GCGS+P 90.73 ±1.63 154 ±9 99.89 ±0.25 20 ±2 66.67 ± 7.86 2984 ±490 DeepSeek Coder 33B GCGS 76.90 ±1.87 283 ±16 95.27 ±0.22 84 ±4 66.67 ± 3.14 3143 ±176 GCGS+P 85.69 ±1.16 240 ±22 96.41 ±0.61 80 ±6 63.97 ± 3.86 3239 ±510 Llama 3.1 8B GCGS 93.73 ±1.57 90 ±9 99.88 ±0.27 22 ±4 48.89 ± 2.48 4059 ±230 GCGS+P 97.11 ±0.66 65 ±8 99.88 ±0.27 22 ±3 54.00 ± 8.94 3719 ±292 Qwen 2.5 Coder 7B GCGS 70.84 ±1.25 317 ±9 81.29 ±1.46 180 ±3 48.33 ± 6.97 3962 ±427 GCGS+P 76.03 ±1.76 299 ±14 84.53 ±1.55 169 ±6 55.00 ± 7.45 3813 ±535 Qwen 2.5 Coder 32B GCGS 43.50 ±1.94 501 ±9 73.17 ±1.68 228 ±8 23.08 ± 5.44 5927 ±328 GCGS+P 50.63 ±1.76 492 ±14 75.37 ±1.48 235 ±7 27.69 ± 4.21 5839 ±281
To evaluate our attack’s effectiveness, we target both closed-source (GPT 4.1 (2025/04/14) [44], Claude 3.5 Sonnet v2 (2024/10/22) [12]) and open-source models (Llama 3.1 8B Instruct [21], Qwen 2.5 Coder Instruct (7B and 32B) [30, 65], DeepSeek Coder Instruct (6.7B and 33B) [27], and Codestral 22B v0.1 [57]). Notably, models like GPT 4.1 and Claude 3.5 Sonnet v2 are currently deployed in production AI assistants such as GitHub Copilot Chat [19]. Without any adversarial transformations, the models achieve an average 68.06% pass@1 rate (see § C.3 for details).
For reproducibility, we set the sampling temperature to 0 for greedy decoding [^2] Anthropic API notes that setting temperature 0.0 does not guarantee complete determinism for its models.. For open-source models, we run our attack with five random seeds. Due to resource constraints, we evaluate closed-source models with a single full run, supplemented by limited runs on small dataset samples using five distinct random seeds (see § H). To demonstrate the efficacy of our perplexity-based deep greedy composition as well as our support for models like Anthropic that do not provide token probabilities, we run our attack both with the token probabilities feedback (GCGS+P) and without (GCGS).
Results. As shown in Table 1, GCGS and GCGS+P demonstrate high effectiveness and efficiency at attacking SoTA LLMs on in-context code generation, with ASR and average number of queries ranging from 40.69% to 99.88% and 22 to 501, respectively. Despite Claude 3.5 Sonnet’s competitive baseline performance, it demonstrates high vulnerability to GCGS, underscoring the attack’s viability without model feedback. However, perplexity-based feedback consistently enhances attack effectiveness, with GCGS+P improving ASR over GCGS by up to 8.79 percentage points, often while requiring fewer queries. The attack proves particularly effective on MBPP+, achieving over 95% ASR on 5 of 8 evaluated models, suggesting this benchmark may be more susceptible to GCGS-style attacks than HumanEval+. Furthermore, within model families, larger variants consistently show greater resilience to the attack (e.g., Qwen 2.5 Coder 32B vs 7B). Although the Qwen 2.5 Coder family shows notably greater resilience than other models on both datasets, GCGS+P still achieves over 50% ASR. GPT 4.1 exhibits anomalous resilience on MBPP+ (40.69% ASR) compared to HumanEval+ (81.82%), but its closed-source nature prevents determining the root cause.
6.2 In-Context Vulnerability Injection
@@
Task Description. To quantify the ability of GCGS to inject vulnerabilities, we evaluate GCGS on the CWEval dataset [46], specifically designed to assess both functionality and security of LLM-generated code. Using our identical EvalPlus setup for § 6.1 on CWEval’s Python subset (CWEval/Python), we measured attack success by target LLMs generating code that passes functional tests and fails security tests linked to specific Common Weakness Enumeration (CWE) categories [6].
Results. Although injecting specific vulnerabilities while preserving preserving functionality is much more difficult than untargeted bug injection, GCGS still successfully triggers 17 unique CWEs across different models, achieving an average ASR of 52.26% (Table 1). Consistent with our results in § 6.1, perplexity-based feedback helps improve performance, with the exception of DeepSeek Coder 33B which we believe might be caused by the dataset’s small size. We’ve added three case studies highlighting the most concerning behaviors observed in the tested models in § I.
6.3 Code Reasoning
@@
Task Description. To evaluate our ability to attack code reasoning LLMs, we select two security-focused binary classification benchmarks from CodeXGLUE [42]: Defect Detection and Clone Detection, both well-established in the adversarial code transformation literature [66, 71, 43]. The Defect Detection task builds on Devign [74], a dataset of 27,318 real-world C functions annotated for security vulnerabilities. The Clone Detection task employs BigCloneBench [56, 61], which includes over 1.7 million labeled code pairs spanning from syntactically identical to semantically similar code fragments. We evaluate our attack on three fine-tuned LLMs that achieve SoTA performance on these tasks: CodeBERT [22], GraphCodeBERT [26], and CodeT5+ 110M [62]. We did not evaluate generative coding models because of their low performance on these tasks (more details in § C.3). We compare against several leading adversarial attacks that leverage semantics-preserving code transformations: ALERT [66] and MHM [69] (chosen for their prevalence in comparative studies), RNNS [71] (a recent performant approach), and WIR-Random [68] (the most effective non-Java-specific attack from a comprehensive study [20]).
@@
Table 2: Performance of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE tasks. Results show mean ± std over 5 seeds. Best ASR per model is in bold.
@@
Defect Detection Clone Detection CodeBERT GraphCodeBERT CodeT5+ CodeBERT GraphCodeBERT CodeT5+ Attack ASR #Queries ASR #Queries ASR #Queries ASR #Queries ASR #Queries ASR #Queries ALERT 62.35 ±5.92 732 ±120 76.87 ±5.00 468 ±106 62.22 ±8.02 784 ±196 19.32 ±6.15 2125 ±161 21.02 ±2.89 2083 ±87 25.35 ±5.16 2008 ±117 MHM 56.48 ±9.14 742 ±111 75.64 ±12.84 479 ±178 82.81 ±1.98 405 ±34 26.10 ±8.98 1000 ±85 32.78 ±6.05 944 ±55 37.97 ±8.25 874 ±74 RNNS 73.97 ±6.39 479 ±67 86.51 ±5.11 331 ±61 86.59 ±3.21 355 ±44 42.87 ±4.27 1036 ±66 44.91 ±4.03 967 ±59 46.90 ±7.59 1045 ±109 WIR-Random 64.82 ±6.65 145 ±11 78.80 ±8.44 125 ±15 74.43 ±2.02 134 ±8 24.76 ±6.48 236 ±15 30.41 ±6.01 224 ±7 31.78 ±6.36 224 ±12 GCGS 93.18 ±5.79 259 ±172 94.11 ±6.13 229 ±155 97.76 ±1.12 177 ±27 72.27 ±5.38 1032 ±106 64.02 ±6.70 1150 ±100 65.34 ±3.82 1078 ±42 GCGS+W 97.17 ±1.90 167 ±45 97.22 ±3.03 147 ±113 99.89 ±0.13 46 ±48 80.97 ±2.48 728 ±113 83.19 ±3.20 545 ±69 69.04 ±8.77 835 ±165
For model training and evaluation, we use different approaches for our two datasets. For Defect Detection, we fine-tune models on the full dataset. For Clone Detection, due to its substantial size, we follow previous literature and use a balanced subset of 90,000 training and 4,000 validation examples to ensure computational feasibility. We sample 400 test examples from Clone Detection to enable multiple evaluations of each attack-model combination. For GCGS’s warm-up, we withhold a small subset of the fine-tuning datasets: 1,000 training and 100 validation examples for Defect Detection, and 4,000 training and 200 validation examples for Clone Detection. The one-time computational costs of warm-up in terms of model queries are detailed in § G. To mitigate the effects of randomness during model fine-tuning and attacking, we fine-tune each model five times on five random seeds and run each attack with the same random seed on each fine-tuned model.
Defect Detection Results. GCGS uses up to 50.14% fewer queries than the next best performer, RNNS, while delivering consistently higher success rates across all evaluated models (Table 2). While WIR-Random achieves lower query counts on CodeBERT and GraphCodeBERT, its success rate falls short of GCGS by a considerable margin of up to 28.36 percentage points. The warmed-up variant (GCGS+W) is particularly performant on CodeT5+, where it approaches perfect attack success while reducing the required queries by 74.01% to just 46 queries on average. Remarkably, GCGS+W achieves this by warming up on just 1,100 examples – a mere 4.02% of the dataset.